# SalUn: Empowering Machine Unlearning via Gradient-based Weight Saliency in Both Image Classification and Generation

### 06 Nov 2023 v2

@@ Abstract
With evolving data regulations, machine unlearning (MU) has become an important tool for fostering trust and safety in today’s AI models. However, existing MU methods focusing on data and/or weight perspectives often grapple with limitations in unlearning accuracy, stability, and cross-domain applicability. To address these challenges, we introduce the concept of ‘weight saliency’ in MU, drawing parallels with input saliency in model explanation. This innovation directs MU’s attention toward specific model weights rather than the entire model, improving effectiveness and efficiency. The resultant method that we call saliency unlearning (SalUn) narrows the performance gap with ‘exact’ unlearning (model retraining from scratch after removing the forgetting dataset). To the best of our knowledge, SalUn is the first principled MU approach adaptable enough to effectively erase the influence of forgetting data, classes, or concepts in both image classification and generation. For example, SalUn yields a stability advantage in high-variance random data forgetting, e.g., with a 0.2% gap compared to exact unlearning on the CIFAR-10 dataset. Moreover, in preventing conditional diffusion models from generating harmful images, SalUn achieves nearly 100% unlearning accuracy, outperforming current state-of-the-art baselines like Erased Stable Diffusion and Forget-Me-Not. Codes are available at  https://github.com/OPTML-Group/Unlearn-Saliency.
@@ 1 Introduction
Machine unlearning (MU) is the problem of efficiently and effectively mitigating the influence of particular data points in a pre-trained model (Shaik et al., 2023). It emerged in response to data protection regulations like ‘the right to be forgotten’ (Hoofnagle et al., 2019). However, its relevance extends to many trustworthy ML problems, such as defending against backdoor poisoning (Liu et al., 2022a), enhancing group fairness (Oesterling et al., 2023), and improving transfer learning (Jia et al., 2023). Roughly speaking, current MU methods can be categorized into two families: exact or certified MU and approximate MU. The former focuses on developing methods with provable error guarantees or unlearning certifications. Examples of such methods include differential privacy (DP)-enforced unlearning and certified data removal (Guo et al., 2019; Chien et al., 2022). Within this family, exact unlearning, which involves retraining a model from scratch after removing the forgetting dataset from the original training set, is typically considered the gold standard of MU (Thudi et al., 2022b, a). However, retraining-based exact unlearning methods require significant computation resources and have become challenging for today’s large-scale ML models, such as the stable diffusion generative model considered in this work.
@@
In contrast to exact or certified MU, approximate unlearning has emerged as a more practical approach for ‘fast’ and ‘accurate’ unlearning. Here, the accuracy may not meet provable guarantees, but it can be evaluated using more diverse and practical metrics, such as membership inference attacks (Carlini et al., 2022), without the need for data-model or algorithmic assumptions that certified unlearning might require. Despite the merits of efficiency and flexibility, the performance of approximate unlearning can still exhibit significant variance. For example, influence unlearning (Izzo et al., 2021; Warnecke et al., 2021), built upon the influence function analysis of training data points (Koh & Liang, 2017), can vary in effectiveness depending on the choice of hyperparameters (needed for influence function approximations) and the specific unlearning scenarios and evaluation metrics (Becker & Liebig, 2022), leading to the instability concern in MU. Other approximate unlearning methods, including Fisher forgetting (Golatkar et al., 2020; Becker & Liebig, 2022), gradient ascent (Thudi et al., 2022a), and finetuning-based approaches (Warnecke et al., 2021; Jia et al., 2023), also face the similar challenge as will be illustrated later.
@@
Furthermore, many MU methods mentioned above have been primarily applied to image classification. By contrast, emerging diffusion models (DMs) for generative modeling also demand effective MU techniques to protect copyrights and prevent generation of harmful content (Schramowski et al., 2023; Gandikota et al., 2023; Zhang et al., 2023). However, as this work will demonstrate, existing MU methods designed for image classification are insufficient to address MU in image generation (see Fig. 1 for a schematic overview of our proposal vs. conventional MU).
@@
Figure 1: Schematic overview of comparing conventional unlearning methods with our proposal (SalUn) in the context of mitigating the influence of the harmful concept of ‘nudity’ in DM.
To tackle (Q), we develop an innovative MU paradigm: ‘weight saliency’. Drawing a parallel with input saliency in model explanation, our idea shifts the spotlight of MU from the entire model to specific, influential weights. Such focused attention can enhance the performance of multiple MU methods, even simple ones such as random labeling. Termed ‘saliency unlearning’ (SalUn), our approach can diminish the performance gap with exact unlearning, offering a principled MU method effective across image classification or generation. Our contributions are as follows. ❶ We identify two limitations of current MU techniques: instability, e.g., when faced with varying amounts of forgetting data, and lack of adaptability to image generation tasks. ❷ We introduce the concept of weight saliency in MU and develop SalUn, a saliency-guided approach. We show that weight saliency could be a key to addressing the limitations of current MU methods. ❸ We perform comprehensive experiments to validate the effectiveness of SalUn, comparing it with 7 MU baselines in image classification and 2 concept-erasing baselines in image generation. As a notable application, we demonstrate that SalUn is the sole effective method in preventing stable diffusion from generating harmful images when given inappropriate prompts (I2P) (Schramowski et al., 2023).
@@ 2 Related Work
Unlearning in image generation. Recent advancements in text-conditioned image generation models have demonstrated remarkable capabilities in producing high-quality images that closely align with textual descriptions (Rombach et al., 2022; Ho & Salimans, 2022). However, these achievements often rely on extensive datasets, such as LAION-400M and LAION-5B (Schuhmann et al., 2021, 2022), which inherently introduce biases and associated risks. These concerns are indicative of broader issues within the field, as highlighted by various studies (Birhane et al., 2021; Schramowski et al., 2023; Somepalli et al., 2023). To address these challenges, there is a pressing need to explore effective MU techniques. While approaches like Erasing Stable Diffusion (ESD) (Gandikota et al., 2023) and Forget-me-not (FMN) (Zhang et al., 2023) offer solutions for concept erasure in diffusion models, the field currently lacks methods specifically tailored for precise unlearning.
Data and model saliency analyses. There has been extensive research on input saliency maps for the development of explainable ML techniques. Examples include pixel-space sensitivity map methods (Simonyan et al., 2013; Zeiler & Fergus, 2014; Springenberg et al., 2014; Smilkov et al., 2017; Sundararajan et al., 2017) and class-discriminative localization methods (Zhou et al., 2016; Selvaraju et al., 2017; Chattopadhay et al., 2018; Petsiuk et al., 2018). In addition, there has also been a growing body of research focused on data-level saliency analyses, often referred to as data attribution (Koh & Liang, 2017; Park et al., 2023; Ilyas et al., 2022). The application of data attribution includes model explanation (Jeyakumar et al., 2020; Grosse et al., 2023), debugging (Ilyas et al., 2022), efficient training (Xie et al., 2023), and improving model generalization (Jain et al., 2023). Compared to input saliency and data attribution, model saliency is a less explored concept. Weight sparsity (Han et al., 2015; Frankle & Carbin, 2018), commonly used in weight pruning to enhance model efficiency, can be viewed as a form of weight saliency map that focuses on preserving a model’s generalization ability. In the field of natural language processing (NLP), research on model editing (Dai et al., 2021; Meng et al., 2022; De Cao et al., 2021) has focused on locating and modifying specific knowledge within a model by directly targeting and modifying model weights. This concept of an ‘editable model region’ aligns with the notion of weight saliency in NLP, where certain model parameters are considered more influential and editable than others.
@@ 3 Preliminaries and Problem Statement
Machine unlearning (MU): Objective and setup. MU has become a vital concept and approach in ML, allowing us to remove the influence of specific data points, data classes, or even higher-level data concepts from a pre-trained ML model without requiring a complete retraining of the model from scratch. The data points earmarked for unlearning are commonly known as the forgetting dataset. Thus, the primary objective of MU can be framed as the efficient and effective updating of a pre-trained ML model, so as to attain performance on par with complete retraining (termed as Retrain) that is achieved after the removal of the forgetting dataset from the training set.
@@
To be concrete, let $\mathcal{D}=\{\mathbf{z}_{i}\}_{i=1}^{N}$ denote the training dataset encompassing $N$ data points (including data feature $\mathbf{x}_{i}$ and label $\mathbf{y}_{i}$ for supervised learning). And let $\mathcal{D_{\mathrm{f}}}\subseteq\mathcal{D}$ be the forgetting dataset. Its complement, denoted by $\mathcal{D_{\mathrm{r}}}=\mathcal{D}\setminus\mathcal{D}_{\mathrm{f}}$, is referred to as the remaining dataset. Prior to MU, we denote the original model as ${\bm{\theta}}_{\mathrm{o}}$}, trained on $\mathcal{D}$ using, e.g., empirical risk minimization (ERM). Consistent with existing literature (Thudi et al., 2022a), we regard Retrain as the MU gold standard, which involves training model parameters ($\bm{\theta}$}) from scratch over $\mathcal{D}_{\mathrm{r}}$. Nonetheless, Retrain can be computationally demanding. Hence, the central challenge in MU is to acquire an unlearned model (referred to as ${\bm{\theta}}_{\mathrm{u}}$}) from ${\bm{\theta}}_{\mathrm{o}}$} on $\mathcal{D}_{\mathrm{f}}$ and/or $\mathcal{D_{\mathrm{r}}}$, so that it can serve as an accurate and computationally efficient substitute for Retrain. In what follows, we introduce two MU paradigms that are the primary focus of this work: MU for image classification and MU for image generation.
@@
MU for image classification. This is the most commonly studied MU problem in the literature (Shaik et al., 2023). Depending on the composition of the forgetting dataset $\mathcal{D}_{\mathrm{f}}$, MU for image classification can be further categorized into two scenarios: class-wise forgetting and random data forgetting. The former aims to eliminate the influence of training data points from the same image class, while the latter aims to remove the influence of randomly selected data points from the entire training set.
@@
Evaluating the effectiveness of MU for image classification has involved the use of various metrics. While a consensus is still lacking, we adhere to the recent approach proposed by (Jia et al., 2023), which considers a comprehensive ‘full-stack’ MU evaluation. This includes unlearning accuracy (UA), (i.e., 1 - accuracy of an unlearned model ${\bm{\theta}}_{\mathrm{u}}$} on $\mathcal{D}_{\mathrm{f}}$), membership inference attack (MIA) on $\mathcal{D}_{\mathrm{f}}$, i.e., the accuracy of leveraging MIA to determine if a forgetting data is not in the training set of ${\bm{\theta}}_{\mathrm{u}}$}, remaining accuracy (RA), i.e., the fidelity of an unlearned model ${\bm{\theta}}_{\mathrm{u}}$} on the remaining training set $\mathcal{D}_{\mathrm{r}}$, testing accuracy (TA), i.e., the generalization of ${\bm{\theta}}_{\mathrm{u}}$}, and run-time efficiency (RTE), i.e., the computation time of applying an MU method. Except for RTE, a good unlearning method is expected to achieve performance close to that of Retrain.
@@
where $\hat{\epsilon}(\mathbf{x}_{t}|c)$ stands for the ultimate noise estimation attained by utilizing the conditional DM given $c$, $w\in[0,1]$ is a guidance weight, and ${\epsilon}(\mathbf{x}_{t}|\emptyset)$ signifies the corresponding unconditional employment of the DM. The inference stage initiates with Gaussian noise $z_{T}\sim\mathcal{N}(0,1)$, which is then denoised using $\hat{\epsilon}_{{\bm{\theta}}}(\mathbf{x}_{T}|c)$ to obtain $z_{T-1}$. This procedure is repeated to generate the authentic data at $t=0$. When training the DM $\bm{\theta}$}, the mean-squared-error (MSE) loss is defined as follows
@@
Given a well-trained DM $\bm{\theta}$}, the objective of MU for image generation is twofold: firstly, to prevent $\bm{\theta}$} from generating undesired image content, e.g., when conditioned on harmful concepts like nudity; secondly, to ensure that the post-unlearning updated DM maintains the quality of image generation for normal images. Finally, it is worth noting that in existing literature, the problem of MU for image generation was not studied through the lens of MU. Instead, it was initially termed as ‘learning to forget’ or ‘concept erasing’ (Schramowski et al., 2023; Gandikota et al., 2023; Zhang et al., 2023). However, we will show that MU provides a systematic framework for addressing this challenge.
@@ 4 Challenges in Current Machine Unlearning Methods
Figure 2: Demonstration of the instability limitations of MU methods on CIFAR-10. (a) Sensitivity of unlearning accuracy gaps with respect to Retrain (measured by $|\text{Method}-\text{{\text{Retrain}}}|$’) as a function of forgetting data amount. Five existing MU methods (FT, RL, GA, IU, $\ell_{1}$-sparse) are included. (b) Box plots illustrating the variances of unlearning accuracies using Retrain, IU, and the proposed weight saliency-integrated IU across various hyperparameter choices. The box size represents the variance of UA values across hyperparameter values.
@@
The instability limitation. In evaluating the performance of MU methods, previous research has often assumed a fixed number of forgetting data, such as data points within an entire class or a fixed fraction of the training set. There has been limited evaluation exploring how the unlearning performance is affected by varying quantities of forgetting data. In Fig. 2(a), we investigate the unlearning performance gap relative to the gold standard Retrain, measured in terms of UA (unlearning accuracy), as a function of the quantity of forgetting data points. Note that a smaller performance gap is desirable. As we can see, the unlearning effectiveness of MU methods (①-⑤) observed at a 10% forgetting data quantity does not necessarily hold when the forgetting data quantity is increased to 50%. Similar instability in other performance metrics can also be observed in more experiment results later.

### 19 Feb 2024 v3

@@ Abstract
With evolving data regulations, machine unlearning (MU) has become an important tool for fostering trust and safety in today’s AI models. However, existing MU methods focusing on data and/or weight perspectives often suffer limitations in unlearning accuracy, stability, and cross-domain applicability. To address these challenges, we introduce the concept of ‘weight saliency’ for MU, drawing parallels with input saliency in model explanation. This innovation directs MU’s attention toward specific model weights rather than the entire model, improving effectiveness and efficiency. The resultant method that we call saliency unlearning (SalUn) narrows the performance gap with ‘exact’ unlearning (model retraining from scratch after removing the forgetting data points). To the best of our knowledge, SalUn is the first principled MU approach that can effectively erase the influence of forgetting data, classes, or concepts in both image classification and generation tasks. For example, SalUn yields a stability advantage in high-variance random data forgetting, e.g., with a 0.2% gap compared to exact unlearning on the CIFAR-10 dataset. Moreover, in preventing conditional diffusion models from generating harmful images, SalUn achieves nearly 100% unlearning accuracy, outperforming current state-of-the-art baselines like Erased Stable Diffusion and Forget-Me-Not. Codes are available at  https://github.com/OPTML-Group/Unlearn-Saliency.
@@ 1 Introduction
Machine unlearning (MU) is the problem of efficiently and effectively mitigating the influence of particular data points on a pre-trained model (Shaik et al., 2023). It emerged in response to data protection regulations like ‘the right to be forgotten’ (Hoofnagle et al., 2019). However, its scope and significance rapidly expand to tackle many trustworthy machine learning (ML) challenges in computer vision (CV). These challenges include the defense against backdoor poisoning attacks (Liu et al., 2022a), the enhancement of model fairness (Oesterling et al., 2023), the refinement of pre-training methods to augment transfer learning capabilities (Jain et al., 2023; Jia et al., 2023), and the prevention of text-to-image generative models from generating sensitive, harmful, or illegal image content when exposed to inappropriate prompts (Gandikota et al., 2023).
@@
Roughly speaking, current MU methods can be categorized into two families: exact or certified MU and approximate MU. The former focuses on developing methods with provable error guarantees or unlearning certifications. Examples of such methods include differential privacy (DP)-enforced unlearning and certified data removal (Guo et al., 2019; Chien et al., 2022). Within this family, exact unlearning, which involves retraining a model from scratch after removing the forgetting dataset from the original training set, is typically considered the gold standard of MU (Thudi et al., 2022b; a). However, retraining-based exact unlearning methods require significant computation resources and have become challenging for today’s large-scale ML models, such as the diffusion-based generative model considered in this work.
@@
In contrast to exact or certified MU, approximate unlearning has emerged as a more practical approach for ‘fast’ and ‘accurate’ unlearning. While the accuracy may not meet provable guarantees, it can b assessed using a broader range of practical metrics, such as membership inference attacks (Carlini et al., 2022), without necessitating data-model or algorithmic assumptions typically associated with certified unlearning. Despite the merits of practicality and efficiency, the performance of approximate unlearning can still exhibit significant variance. For example, influence unlearning (Izzo et al., 2021; Warnecke et al., 2021), built upon the influence function analysis of training data points (Koh & Liang, 2017), exhibits high performance variance due to the selection of hyperparameters required for influence function approximations, as well as the particular unlearning scenarios and evaluation metrics (Becker & Liebig, 2022), thereby raising concerns about instability in approximate unlearning methods. Other approximate unlearning methods, including Fisher forgetting (Golatkar et al., 2020), gradient ascent (Thudi et al., 2022a), and finetuning-based approaches (Warnecke et al., 2021; Jia et al., 2023), also face the similar challenge as will be illustrated later.
@@
Furthermore, many MU methods mentioned above have been primarily applied to image classification. By contrast, emerging diffusion models (DMs) for generative modeling also demand effective MU techniques to protect copyrights and prevent generation of harmful content (Schramowski et al., 2023; Gandikota et al., 2023; Zhang et al., 2023a). However, as this work will demonstrate, existing MU methods designed for image classification are insufficient to address MU in image generation (see Fig. 1 for a schematic overview of our proposal vs. conventional MU).
@@
Figure 1: Schematic overview of our proposal (SalUn) vs. the conventional unlearning method in the context of removing the influence of the harmful concept ‘nudity’ in diffusion generation.
To tackle (Q), we develop an innovative MU paradigm: ‘weight saliency’. Drawing a parallel with input saliency in model explanation, our idea shifts the spotlight of MU from the entire model to specific, influential weights. Such focused attention can enhance the performance of multiple MU methods, even simple ones such as random labeling. Termed ‘saliency unlearning’ (SalUn), our approach can diminish the performance gap with exact unlearning, offering a principled MU method effective across image classification or generation. Our contributions are as follows. ❶ We identify two limitations of current MU techniques: instability, e.g., when faced with varying amounts of forgetting data, and lack of adaptability to image generation tasks. ❷ We introduce the concept of weight saliency in MU and develop SalUn, a saliency-guided approach. We show that weight saliency could be a key to addressing the limitations of current MU methods. ❸ We perform comprehensive experiments to validate the effectiveness of SalUn, comparing it with 7 MU baselines in image classification and 2 concept-erasing baselines in image generation. As a notable application, we show that SalUn is the most effective method in preventing stable diffusion from generating harmful images when given inappropriate prompts (I2P) (Schramowski et al., 2023).
@@ 2 Related Work
Unlearning in image generation. Recent advancements in text-conditioned image generation models have demonstrated remarkable capabilities in producing high-quality images that closely align with textual descriptions (Rombach et al., 2022; Ho & Salimans, 2022). However, these achievements often rely on extensive datasets, such as LAION-400M and LAION-5B (Schuhmann et al., 2021; 2022), which inherently introduce biases and associated risks. These concerns are indicative of broader issues within the field, as highlighted by various studies (Birhane et al., 2021; Schramowski et al., 2023; Somepalli et al., 2023; Bae et al., 2023; Zhang et al., 2023b). To address these challenges, there is a pressing need to explore effective MU techniques. While current studies (Gandikota et al., 2023; Zhang et al., 2023a; Heng & Soh, 2023) provide strategies for concept erasure in diffusion models, achieving precision comparable to exact unlearning remains challenging.
Data and model saliency analyses. There has been extensive research on input saliency maps for the development of explainable ML techniques. Examples include pixel-space sensitivity map methods (Simonyan et al., 2013; Zeiler & Fergus, 2014; Springenberg et al., 2014; Smilkov et al., 2017; Sundararajan et al., 2017) and class-discriminative localization methods (Zhou et al., 2016; Selvaraju et al., 2017; Chattopadhay et al., 2018; Petsiuk et al., 2018). In addition, there has also been a growing body of research focused on data-level saliency analyses, often referred to as data attribution (Koh & Liang, 2017; Park et al., 2023; Ilyas et al., 2022). The application of data attribution includes model explanation (Jeyakumar et al., 2020; Grosse et al., 2023), debugging (Ilyas et al., 2022), efficient training (Xie et al., 2023), and improving model generalization (Jain et al., 2023). Compared to input saliency and data attribution, model saliency is a less explored concept. Weight sparsity (Han et al., 2015; Frankle & Carbin, 2018), commonly used in weight pruning to enhance model efficiency, can be viewed as a form of weight saliency map that focuses on preserving a model’s generalization ability. In the field of natural language processing (NLP), research on model editing (Dai et al., 2021; Meng et al., 2022; De Cao et al., 2021; Patil et al., 2023) has focused on locating and modifying specific knowledge within a model by directly targeting and modifying model weights. This concept of an ‘editable model region’ aligns with the notion of weight saliency in NLP, where certain model parameters are considered more influential and editable than others.
@@ 3 Preliminaries and Problem Statement
Machine unlearning (MU): Objective and setup. MU has become a vital concept and approach in ML, allowing us to remove the influence of specific data points, data classes, or even higher-level data concepts from a pre-trained ML model without requiring a complete retraining of the model from scratch. The set of data points earmarked for unlearning is commonly known as the forgetting dataset. Thus, the primary objective of MU can be framed as the efficient and effective updating of a pre-trained ML model, so as to attain performance on par with complete retraining (termed as Retrain), which is achieved after the removal of the forgetting dataset from the training set.
@@
To be concrete, let $\mathcal{D}=\{\mathbf{z}_{i}\}_{i=1}^{N}$ denote the training dataset encompassing NN data points (including data feature $\mathbf{x}_{i}$ and label $\mathbf{y}_{i}$ for supervised learning). And let $\mathcal{D_{\mathrm{f}}}\subseteq\mathcal{D}$ be the forgetting dataset. Its complement, denoted by $\mathcal{D_{\mathrm{r}}}=\mathcal{D}\setminus\mathcal{D}_{\mathrm{f}}$, is referred to as the remaining dataset. Prior to MU, we denote the original model as ${\bm{\theta}}_{\mathrm{o}}$}, trained on $\mathcal{D}$ using, e.g., empirical risk minimization (ERM). Consistent with existing literature (Thudi et al., 2022a; Jia et al., 2023), we regard Retrain as the MU’s gold standard, which involves training mode parameters ($\bm{\theta}$}) from scratch over $\mathcal{D}_{\mathrm{r}}$. Nonetheless, Retrain can be computationally demanding. Hence, the central challenge in MU is to acquire an unlearned model (referred to as ${\bm{\theta}}_{\mathrm{u}}$}) from ${\bm{\theta}}_{\mathrm{o}}$} on $\mathcal{D}_{\mathrm{f}}$ and/or $\mathcal{D_{\mathrm{r}}}$, so that it can serve as an accurate and computationally efficient substitute for Retrain. In what follows, we introduce two MU paradigms that are the primary focus of this work: MU for image classification and MU for image generation.
@@
MU for image classification. This is the most commonly studied MU problem in the literature (Shaik et al., 2023). Depending on the composition of the forgetting dataset $\mathcal{D}_{\mathrm{f}}$, MU for image classification can be further categorized into two scenarios: class-wise forgetting and random data forgetting. The former aims to eliminate the influence of an image class, while the latter aims to remove the influence of randomly selected data points from the entire training set.
@@
Evaluating the effectiveness of MU for image classification has involved the use of various metrics. While a consensus is still lacking, we adhere to the recent approach proposed by (Jia et al., 2023), which considers a comprehensive ‘full-stack’ MU evaluation. This includes unlearning accuracy (UA), i.e., 11 - the accuracy of an unlearned model ${\bm{\theta}}_{\mathrm{u}}$} on $\mathcal{D}_{\mathrm{f}}$, membership inference attack (MIA) on $\mathcal{D}_{\mathrm{f}}$, i.e., the privacy measure of ${\bm{\theta}}_{\mathrm{u}}$} over $\mathcal{D}_{\mathrm{f}}$, remaining accuracy (RA), i.e., the fidelity of an unlearned model ${\bm{\theta}}_{\mathrm{u}}$} on the remaining training set $\mathcal{D}_{\mathrm{r}}$, testing accuracy (TA), i.e., the generalization of ${\bm{\theta}}_{\mathrm{u}}$}, and run-time efficiency (RTE), i.e., the computation time of applying an MU method.
@@
where $\hat{\epsilon}(\mathbf{x}_{t}|c)$ stands for the ultimate noise estimation attained by utilizing the conditional DM given $c$, $w\in[0,1]$ is a guidance weight, and ${\epsilon}(\mathbf{x}_{t}|\emptyset)$ signifies the corresponding unconditional employment of the DM. The inference stage initiates with Gaussian noise $z_{T}\sim\mathcal{N}(0,1)$, which is then denoised using $\hat{\epsilon}_{{\bm{\theta}}}(\mathbf{x}_{T}|c)$ to obtain $z_{T-1}$. This procedure is repeated to generate the authentic data at $t=0$. When training the DM $\bm{\theta}$}, the mean-squared-error (MSE) loss is commonly used
@@
Given a well-trained DM $\bm{\theta}$}, the objective of MU for image generation is twofold: (1)preventing $\bm{\theta}$} from generating undesired image content, e.g., when conditioned on harmful concepts like nudity, and (2) ensuring that the post-unlearning updated DM maintains the quality of image generation for normal images. Finally, it is worth noting that in existing literature, the problem of MU for image generation was not studied through the lens of MU. Instead, it was initially termed as ‘learning to forget’ or ‘concept erasing’ (Schramowski et al., 2023; Gandikota et al., 2023; Zhang et al., 2023a). However, we will show that MU provides a systematic framework for addressing this challenge.
@@ 4 Challenges in Current Machine Unlearning Methods
Figure 2: The instability limitations of MU methods on CIFAR-10. (a) Sensitivity of unlearning accuracy gaps with respect to Retrain (measured by $|\text{Method}-\text{{\text{Retrain}}}|$’) as a function of forgetting data amount. Four MU methods (FT, RL, IU, $\ell_{1}$-sparse) are included. (b) Box plots illustrating unlearning accuracy using Retrain, IU, and the proposed weight saliency-integrated IU across various hyperparameter choices. The box size represents the variance of UA against hyperparameter values.
@@
The instability limitation. In evaluating the performance of MU methods, previous research has often assumed a fixed number of forgetting data, such as data points within an entire class or a fixed ratio of the training set. There has been limited evaluation exploring how the unlearning performance is affected by varying quantities of forgetting data. In Fig. 2(a), we investigate the unlearning performance gap relative to the gold standard Retrain, measured in terms of UA (unlearning accuracy), as a function of the quantity of forgetting data points. Note that a smaller performance gap is desirable. As we can see, the unlearning effectiveness of MU methods (①-⑤) observed at a 10% forgetting data quantity does not necessarily hold when the forgetting data quantity is increased to 50%. Similar instability in other performance metrics can also be observed in more experiment results later.